Mochi:

Anthony is co-hosting a graduation party. He asked the guests who will attend to select cake or pie for dessert. Cake was selected by 98 guests, and pie was selected by 67 guests. If cakes are cut into 12 slices each, and pies are cut into 8 slices each, how many cakes and how many pies does Anthony need to bake? A. 7 pies and 6 cakes B. 8 cakes and 10 pies C. 9 cakes and 9 pies D. 8 cakes and 8 pies

IsabelFriedmannbells:

i think its letter c

jhonyy9:

cake was selected by 98 guests pie was selected by 67 guests cake cuts in 12 slices pie into 8 slices use the rule of 3 simple and write 1 cake give 12 slices x cake give 98 slices cross multiplie and calcule the x value x = 98*1/12 = 98/12 = 8,1 => 9 cakes need to bake 1 pie give 8 slices x pie give 67 slices x = 67*1/8 = 67/8 = 8,3 => 9 pie need to bake hope helped understandably easy

jhonyy9:

questions ? @Laylalyssa